Lorna is twice as older as peter if their ages add up to 9 years old,how old is peter
Found 2 solutions by Zucchini, macston:
Answer by Zucchini(70)

Peter: x
Lorna: 2x
2x + x = 9
3x = 9
x= 3
Peter is 3 years old.

Answer by macston(5194)

P=peter's age; L=Lorna's age=2P
L+P=9 years Substitute for L
2P+P=9 years
3P=9years Divide each side by 3
P=3 years ANSWER 1: Peter age 3 years.
L=2P=2(3 years)=6 years ANSWER 2 : Lorna's age is 6 years.
CHECK
Sum of ages is 9 years.
L+P=9 years
6 years+ 3 years=9 years
9 years=9 years